OpenStudy (chmvijay):

In toluene alkylation Toleune + CH3Cl- > o-xylene +p-xylene + m-xylene which is predominant isomerformed in this reaction ? and why?

OpenStudy (aaronq):

you would get both the ortho and para positions because alkyl groups are o-,p-directing. Alkyl groups promote substitution at those places because of induction and resonance.

OpenStudy (somy):

T_T im studying this resonance thingy with Kai and i still don't get it (btw @Kainui *happy* now i know this reactions also hehe, not resonance tho T_T) @aaronq do u know any source that has explanation oh how to know which is orth/para or meta directors? ~_~ meaning about resonance thing

OpenStudy (aaronq):

hm you could start here: https://www.youtube.com/watch?v=oeX0P1XIEuY You should try to draw the resonance structures of the ring on your own too.

OpenStudy (aaronq):

This one also https://www.youtube.com/watch?v=E9vc9Zmgd1g

OpenStudy (somy):

this is whats the problem >_< okay i'll check them out thnx :*

OpenStudy (aaronq):

okay coool

OpenStudy (somy):

do Deactivators are the ones that take electron and Activators are the ones that donate electron Electron donating groups are ortho/para directors and weakly deactivating halogen is also ortho/para director Electron taking groups are meta directors/ anything that is more deactivating than halogens are meta directors okay i get this one now question CH3 that attached to benzene ring where is the electron that is being donated? or wait groups like CH3 donate electrons TO electophiles and groups like NO2 donate their lone pair to the ring itself!!! hmmm so when the electron from the group is donated to electrophile- it'll be ortho/para director and activator and if an electron from the group is donated to the ring itself then its meta director and deactivator alright idk if im overdoing it ~_~

OpenStudy (aaronq):

The CH3 is donating electron density from the C-H \(\sigma\)-bond to the \(\pi\)-system this increases it's nucleophilicity (and ability to attack electrophiles)  It sounds like you have the right idea though

OpenStudy (somy):

hmmm so its given to pi bond and from there its given to electrophile?

OpenStudy (aaronq):

yeah, the electrons from the \(\pi\)-system form the bond to the electrophile.

OpenStudy (aaronq):

recall they're not "given", but merely shared

OpenStudy (somy):

i see (yeah i guessed it'd be shared also since there will be a bond between electrophile and that particular area in the ring) now about NO2 lone pair is donated to the ring itself this is what i said now i feel like its not really right

OpenStudy (aaronq):

The NO2 group pulls electrons from the \(\pi\)-system,  Take a look at this page and draw the structures on the second picture http://chemwiki.ucdavis.edu/Organic_Chemistry/Hydrocarbons/Aromatics/Properties_of_Aromatics/Inductive_Effects_of_Alkyl_Groups